# You Can Wash Hands Better: Accurate Daily Handwashing Assessment with Smartwatches

### v3

@@ 1 Introduction
We apply the bottom-up schema in UWash for it bypasses segmenting problem in the top-down schema. To tackle the intrinsic false classification of the bottom-up schema, we no longer conduct gesture recognition over the window, but conduct gesture semantic segmentation over the window. Semantic segmentation in the computer vision community is to align every pixel in an image with an object category, aka pixel-wise classification. In UWash, we adapt U-Net [28], a well-known semantic segmentation network for medical images, to align every sampling point (sample) in a sliding window with a gesture category. As the window slides, every sample in the time-serial IMU readings will have gesture alignment for itself, aka sample-wise gesture classification. To facilitate the gesture semantic segmentation results, Pyramid Pooling Module [29] and channel-wise attention [30] are applied to the U-Net. Fig. 1 shows an example of gesture semantic segmentation on a handwashing procedure. With the estimated procedure segmentation, UWash is able to detect the start and end of handwashing, estimate the duration of each gesture, and score each gesture as well as the entire procedure according to the estimated duration.
@@ 4.2 Post Smoothing Methods
TABLE II: Professional Duration of Each Gesture. We select 12 videos to record the recommended duration of each gesture in these videos. Removing the maximum and minimum of each gesture, we take the averages as the professional handwashing duration.
@@ 4.3 Handwashing Scoring
Table II shows the recommended duration of each gesture in these videos. In addition, we remove the maximum and minimum of each gesture and compute the average as the professional handwashing duration, denoted as $D_{i}^{p},i\in[1,2,...,9]$.
We have two empirical assumptions. (1) Since each gesture emphasizes cleaning one part of the hands, we assume each gesture is equally important in handwashing. (2) The quality of cleaning of each gesture increases linearly with its duration, and the perfect quality is reached and saturated when the duration is equal to or greater than the professional duration. Given these assumptions, we score handwashing with Equation 2.
@@
where $\frac{100}{9}$ is the best score of each gesture, to match the first assumption; $D_{i}^{e}$ represents the estimated duration of the $i$-th gesture; $min(1,\frac{D_{i}^{e}}{D_{i}^{p}})$ matches the second assumption. Finally, the $Score$ of a handwashing procedure is the sum of estimated scores of all nine gestures.
@@ 5.4 Cross-Domain Results
As shown in Fig. 12, almost all not-trained participants have low handwashing scores, indicating that they do not adhere to the WHO-recommended handwashing procedure in their daily life. After training, their handwashing scores significantly improve, reaching an average score of approximately 85. When participants deliberately shuffle the handwashing steps or omit steps, UWash can detect a decline in their scores, demonstrating its effectiveness in identifying out-of-sequence guestures or missed steps. In practical use, UWash can alert users to any omitted steps, thereby enhancing handwashing quality in a targeted manner.
@@ 6 Conclusion
In this paper, we introduce UWash, a smartwatch-only handwashing assessment system, to raise people’s awareness of handwashing in daily use and adherence to the WHO handwashing guidelines. UWash takes the data of the IMU sensors of smartwatches, i.e., accelerometers and gyroscopes, as inputs, feeds the inputs into a dual-branch U-Nets, and outputs sample-wise gesture recognition results effectively. UWash can detect handwashing start/end time, estimate the duration of every handwashing gesture, and score gestures as well as the entire procedure following WHO guidelines. Experimental results over 51 participants show that UWash works well in the user-dependent, cross-participant, and cross-participant-cross-location settings. In addition, UWash still performs promising 9 months later in a hospital and 2 years later in a hall during an in-wild test. Moreover, UWash models are lightweight, only 496KB, with great potential to be deployed on edge devices in the future.

### v4

# You Can Wash Hands Better: Accurate Daily Handwashing Assessment with a Smartwatch

@@ 1 Introduction
We apply the bottom-up schema in UWash for it bypasses segmenting problem in the top-down schema. To tackle the intrinsic false classification of the bottom-up schema, we no longer conduct gesture recognition over the window, but conduct gesture semantic segmentation over the window. Semantic segmentation in the computer vision community is to align every pixel in an image with an object category, aka pixel-wise classification. In UWash, we adapt U-Net [28, 29], a well-known semantic segmentation network for medical images, to align every sampling point (sample) in a sliding window with a gesture category. As the window slides, every sample in the time-serial IMU readings will have gesture alignment for itself, aka sample-wise gesture classification. To facilitate the gesture semantic segmentation results, Pyramid Pooling Module [30] and channel-wise attention [31] are applied to the U-Net. Fig. 1 shows an example of gesture semantic segmentation on a handwashing procedure. With the estimated procedure segmentation, UWash is able to detect the start and end of handwashing, estimate the duration of each gesture, and score each gesture as well as the entire procedure according to the estimated duration.
@@ 4.2 Post Smoothing Methods
TABLE II: Professional Duration of Each Gesture. We select 12 videos to record the recommended duration of each gesture in these videos. For each gesture, the maximums (highlighted in red) and minimums (highlighted in blue) are excluded to mitigate outlier effects. The final reference duration for each gesture is the average of the remaining 10 videos.
@@ 4.3 Handwashing Scoring
Table II shows the recommended duration of each gesture in these videos. In addition, we remove the maximum and minimum of each gesture and compute the average as the professional handwashing duration, denoted as $D_{i}^{p},i\in[1,2,...,9]$. We have two empirical assumptions. (1) Since each gesture emphasizes cleaning one part of the hands, we assume each gesture is equally important in handwashing. (2) We assume the quality of cleaning of each gesture increases linearly with its duration, and the perfect quality is reached and saturated when the duration is equal to or greater than the professional duration. Given these assumptions, we score handwashing with Equation 2.
@@
where $\frac{100}{9}$ is the best score of each gesture, to match the first assumption; $D_{i}^{e}$ represents the estimated duration of the $i$-th gesture; $min(1,\frac{D_{i}^{e}}{D_{i}^{p}})$ matches the second assumption. Finally, the $Score$ of a handwashing procedure is the sum of estimated scores of all nine gestures. Note that the above two assumptions only consider handwashing duration and do not incorporate critical quality determinants of hygiene efficacy, e.g., soap coverage and scrubbing vigor.
obscure differences in gesture quality (e.g., whether soap was properly applied or hands fully covered).
@@ 5.4 Cross-Domain Results
As shown in Fig. 12, almost all not-trained participants have low handwashing scores, indicating that they do not adhere to the WHO-recommended handwashing procedure in their daily life. After training, their handwashing scores significantly improve, reaching an average score of approximately 85. When participants deliberately shuffle the handwashing steps or omit steps, UWash can detect a decline in their scores, demonstrating its effectiveness in identifying out-of-sequence gestures or missed steps. In practical use, UWash can alert users to any omitted steps, thereby enhancing handwashing quality in a targeted manner.
@@
6 Discussion and Limitation
There are still several aspects of UWash that need further discussion.
(1) Handwashing Correctness: The current UWash system reduces hand hygiene quality assessment to merely the duration of handwashing actions, neglecting critical factors such as thorough coverage of surfaces, correct motions, soap coverage, and scrubbing vigor that significantly impact cleaning efficacy. A potential solution involves leveraging parametric hand pose or mesh estimation techniques to characterize fine-grained hand movements. For instance, employing the MANO model [41] for hand pose estimation allows for more precise modeling of hand postures. Standalone smartwatches may struggle to capture comprehensive hand pose. A recent work, Ring-a-Pose [42], demonstrates high-accuracy hand pose estimation through a smart ring worn on the middle finger. Future iterations of UWash could collaborate with devices like smart rings to acquire hand pose or mesh, thereby enabling more precise hand hygiene quality assessments.
(2) Smartwatch Placement: The current UWash evaluation assumes a standard watch orientation during testing. However, real-world usage introduces variability—such as loose-fitting wear, inner-wrist positioning, or placement on the non-dominant arm—that challenges system robustness. To address this, additional data collection under such conditions could be implemented, though this approach demands significant human labor. A more labor-efficient alternative involves leveraging synthetic data by establishing a hand coordinate system and injecting orientation-specific perturbations into IMU signals to replicate coordinate system instabilities across distinct wearing positions and varying tightness levels. Domain adaptation [43] and domain generalization [44] techniques could also be explored to mitigate these challenges.
TABLE VIII: UWash latency on a 2010-launched Intel Core i7-620M desktop (seconds).
stride clip inference time sequence inference time post-processing time score calculation time total time
1 0.0067±0.0000 20.4945±4.0483 0.3477±0.0680 0.0026±0.0004 20.8448±4.1167
10 0.0067±0.0002 2.0488±0.3920 0.1319±0.0256 0.0025±0.0004 2.1832±0.4177
20 0.0067±0.0003 1.0311±0.2051 0.1207±0.0234 0.0025±0.0004 1.1543±0.2284
30 0.0067±0.0000 0.6789±0.1321 0.1169±0.0225 0.0026±0.0004 0.7984±0.1550
40 0.0068±0.0003 0.5185±0.0983 0.1141±0.0222 0.0025±0.0004 0.6351±0.1205
50 0.0067±0.0002 0.4131±0.0811 0.1139±0.0223 0.0025±0.0004 0.5295±0.1035
64 0.0067±0.0000 0.3207±0.0619 0.1113±0.0217 0.0025±0.0004 0.4345±0.0840
(3) Power Consumption and Latency: UWash has not been deployed on smartwatches for real-world evaluation of battery consumption and inference latency, despite its parameter count being only 3.2% of MobileNetV3-small [39], a classic deep network optimized for edge devices. In practical deployment, the UWash system architecture can be designed as follows: On the smartwatch end, IMU sensor data is sampled and transmitted to a paired smartphone via Bluetooth Low Energy (BLE). The IMU module, operating at 5 mW (typical for MEMS sensors in continuous sampling mode), and the BLE transmission, consuming 8 mW under a 1 Mbps data rate with 1 ms connection intervals, combine for a total power draw of 13 mW, enabling a smartwatch with a 300 mAh battery capacity (e.g., Samsung Gear Sport) to achieve approximately 115.4 hours (5V × 0.3Ah/0.013W $\approx$ 115.4h) of continuous operation. Meanwhile, the smartphone handles real-time handwashing assessment using the PyTorch Mobile inference engine and relays the results back to the smartwatch via BLE for user feedback.
We evaluated UWash on a desktop-class CPU, the i7-629M (released in 2010), simulating 50 iterations of 60-second handwashing sequences. We measured the average network inference time, post-processing time, and computational steps for scoring, with results summarized in Table VIII. As shown in Table VIII, with a sliding window stride of one sample point during inference, the total processing time was 20.8448±4.1167 seconds, dominated by the sequence inference time of 20.4945±4.0483 seconds. When the stride was increased to 10, the total time decreased to 2.1832±0.4177 seconds, demonstrating that UWash can process a 60-second sequence in approximately 2.1832±0.4177 seconds. For clip-level inference, UWash processes a 1.28-second IMU clip in just 0.0067 seconds.
According to benchmarking data from https://www.cpubenchmark.net/cpu_list.php, the Intel Core i7-629M performs comparably to mobile chips like the Qualcomm Snapdragon 835 and Apple A9. Representative devices using these chips include the 9-year-old Samsung Galaxy S8 and Xiaomi 6 (Snapdragon 835), and the 11-year-old Apple iPhone 6S (A9). These results give a rough estimate of feasibility that mainstream smartphones are capable of supporting real-time execution of UWash, but actual deployment performance may vary due to real-world conditions.
(4) Other Brands of Smartwatches: To evaluate the UWash on other brands of smartwatches, we further recruited 10 new subjects and instructed them to perform five handwashing sessions while wearing the Apple Watch S5. Due to discrepancies in the IMU sensor coordinate systems between the Apple Watch S5 and Samsung Gear Sport, combined with the lack of publicly available parameters for their respective coordinate frameworks, direct parameter alignment remains a significant challenge. Nevertheless, through simple axis swapping—specifically interchanging the x- and y-axes in the Apple Watch data—we tested the Samsung-trained model on Apple Watch data, achieving cross-device performance metrics of 76.24% accuracy, 0.39s start error, 0.57s end error, and 18.81-point scoring error. These results highlight the generalizability of the UWash framework across heterogeneous devices. To improve alignment accuracy, future implementations would require obtaining precise transformation matrices, including the rotation matrix and translation parameters.
Different smartwatch models may exhibit varying IMU sampling rates. For devices with sampling rates below 50 Hz (the default sampling rate of UWash), upsampling techniques such as linear interpolation could be applied to achieve the target 50 Hz frequency prior to model inference. However, this compensatory strategy inherently fails to recover critical kinematic details lost during rapid hand motion transitions, which clinically manifests as reduced detection fidelity in hand hygiene monitoring. We also could downsample the training data to match the target device’s native sampling rate during model training. For devices with sampling rates exceeding 50 Hz, raw data can be downsampled to 50 Hz to align with UWash.
(5) Border Impacts: Our proposed framework is a general-purpose foundation for action recognition and segmentation, with applicability extending far beyond the domain of handwashing. By integrating agent technologies and advanced reasoning capabilities, the system can evolve beyond passive monitoring to deliver actionable insights. For instance, it could generate behavioral summaries to track long-term skill development, provide context-aware feedback tailored to individual user patterns [45], or offer real-time corrective guidance during task execution [46]. This integration not only improves the system’s utility but also opens avenues for personalized skill augmentation, adaptive training interfaces, and human-AI collaboration in complex workflows.
@@ 7 Conclusion
In this paper, we introduce UWash, a smartwatch-only handwashing assessment system, to raise people’s awareness of handwashing in daily use and adherence to the WHO handwashing guidelines. UWash takes the data of the IMU sensors of smartwatches, i.e., accelerometers and gyroscopes, as inputs, feeds the inputs into dual-branch U-Nets, and outputs sample-wise gesture recognition results effectively. UWash can detect handwashing start/end time, estimate the duration of every handwashing gesture, and score gestures as well as the entire procedure following WHO guidelines. Experimental results over 51 participants show that UWash works well in the user-dependent, cross-participant, and cross-participant-cross-location settings. In addition, UWash still performs promising 9 months later in a hospital and 2 years later in a hall during an in-wild test. Moreover, UWash models are lightweight, only 496KB, with great potential to be deployed on edge devices in the future.